OpenStudy (anonymous):

You start a bank account with $50. Each week after that, you deposit $12. How many weeks will it take to have $230 in the account

OpenStudy (anonymous):

You like the algebra way or the common sense way?

OpenStudy (anonymous):

Tell me which way you want to go with this: Algebra or Common Sense Up to you....

OpenStudy (anonymous):

50 + 12x = 230 (x being # of weeks) Now solve for x. Can you do this or do you need help ?

OpenStudy (anonymous):

Or common sense. How much money do we need to deposit to increase the balance from 50 to 230 ? 180, right? How many deposits of $12 will result in $180? 12 times what is 180?

OpenStudy (anonymous):

12*15=180 so 15 weeks @sns6320